OpenStudy (anonymous):

Which of the graphs below represent the graphical solution to the following system of equations? 4x + y = 5 3x – 7y = 27

OpenStudy (anonymous):

you can multiply both 3 and for to elimininate the x in equation; 3/4x+y=5 -4/3x-7y=27

OpenStudy (anonymous):

and then

OpenStudy (anonymous):